# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v2

@@ 5 Conclusion
Ensemble Kalman inversion (EKI) and its variants have proven versatile and easily implemented to efficiently solve optimization problems with random loss functions. Such behavior arises in learning algorithmic parameters in randomized algorithms, e.g., those found in randomized machine learning approaches. This investigation explored a concrete implementation applied to random features regression for model emulation. The main developments in this work were threefold. First, by viewing random features emulators as Gaussian processes, the paper used empirical Bayesian arguments to construct a cost function for hyperparameter estimation. Second, this cost function was adapted to an inverse problem framework compatible with EKI. Third, the feature distributions trained with EKI produced emulators that perform comparably to Gaussian processes and display robustness to overfitting and feature selection.
@@
When trained, random features can offer approximation with fewer assumptions and computational requirements than Gaussian processes. As illustrated in Section 4.3, vector-valued random features were able to produce qualitatively similar posterior distributions without relying on diagonalization, though the framework can take advantage of this when known. As with Gaussian processes, the choice of kernel or feature map is important. In this investigation, the random Fourier features parametrized by low-rank perturbations (Subsection 3.4) were effective across all test cases, given a suitably large rank. On the other hand, multi-output kernel structures that separate input and output spaces led to poor results. Further investigation of sensible kernel selection is left to future work.
The hyperparameter training on the outset can appear quite costly. However reduction of training costs can be tackled on three fronts. First, use of the low-rank-perturbed kernel structure of Subsection 3.4 reduced the number of learnable parameters. Though imposing specific structure, we observe in several applications that this family is sufficiently flexible for a range of tasks. Second, newer variants of EKI use sampling error correction methods to decouple the relationship between parameter dimension and ensemble size (Morzfeld et al., 2024). Such scalability properties are inherited from the ensemble Kalman filtering literature, where routinely assimilators must estimate states of size $\mathcal{O}(10^{9})$ with an ensemble size of $\mathcal{O}(10)$. Third, as the training framework learns the hyperparameters of the feature distribution, the number of samples (features) is a free parameter, and in practice, is often smaller for hyperparameter learning (where accuracy is less important) and larger for prediction tasks (where accuracy is the target). To explore the practical significance of such benefits, comparison with other fixed-feature approaches is an interesting direction for further work.
@@
A fruitful area for future development for this case study is in alternative choices of loss function. More rigorous treatment of this approach may improve optimizer performance and accelerate convergence. Alternative loss functions not motivated from Bayesian arguments also exist for kernel methods and are based on cross validation (Owhadi and Yoo, 2019; Chen et al., 2021; Naslidnyk et al., 2023); such loss functions have proven successful in derivative-based settings for deterministic kernels and could also be recast for random objective functions in the present derivative-free framework. Furthermore, as EKI solves an inverse problem, misfits appearing in the cost function need not be directly supervised, and so this opens the door to training hyperparameters from indirect data. Indirect data is important for many applications where one embeds a machine learning surrogate in place of parametrization or model error compensation, and requires online training (e.g., Pahlavan et al., 2024; Wu et al., 2024).
@@
More broadly, this work intends to act as a case study for a general framework for learning distributions that are sampled in randomized algorithms. Nearby examples are other machine learning tools that follow a randomization approach (Scardapane and Wang, 2017), such as Nyström methods (Williams and Seeger, 2000; Sun et al., 2015; Meanti et al., 2022) or reservoir computers (Gauthier, 2018; Gauthier et al., 2021; Griffith et al., 2019). Common patterns to tune such methods often follow the literature of random features: employing manual calibration, fitting to one realization of randomness, and bespoke algorithms to solve the problem for a narrow choice of algorithm parameters. Recent work in these fields has seen other Bayesian optimization methods learn low-dimensional (fewer than $10$) hyperparameters effectively (e.g., Griffith et al. (2019); Racca and Magri (2021); Antonik et al. (2023)). As EKI is highly scalable when compared with such Gaussian process-based optimizers, it may enable the learning of much higher dimensional hyperparameters. Our approach lays groundwork for new automated approaches of calibration which could accelerate research, as added robustness may lead to quicker uptake of such approaches.
@@
The code and examples for this work are available in three open-source software packages written in the Julia programming language. RandomFeatures.jl v0.3.3 contains a range of different random feature families, EnsembleKalmanProcesses.jl v1.1.5 contains a range of EKI variants and utilities for creating and solving inverse problems, and CalibrateEmulateSample.jl v0.5.3 contains the code for the application problems presented in this paper.
@@
ORAD is supported by Schmidt Sciences, LLC, the National Science Foundation (NSF) under award number AGS-1835860, the Defense Advanced Research Projects Agency (Agreement No. HR00112290030). NHN acknowledges support from the NSF Graduate Research Fellowship Program under award number DGE-1745301, the Amazon/Caltech AI4Science Fellowship, the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), and the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under Office of Naval Research award number N00014-22-1-2790. The authors are grateful to Daniel J. Gauthier for helpful comments on a previous version of the paper.
@@ Appendix A Additional Numerical Experiment Details
In this appendix, we add some additional notes about the experimental configuration to facilitate reproducibility. Note that all numerical experiments are reproducible at
CalibrateEmulateSample.jl v0.5.3.
Users may interact with this code base as a playground to understand our EKI-based algorithm and how its various setup parameters influence performance.
@@ A.1 Hyperparameter Priors
To apply EKI, priors must be placed on learnable parameters. From the practical perspective, priors serve as a method of introducing information, such as enforcing bounds, setting the initial ensemble span and correlation. In particular, for linear problems, the standard EKI update will be confined to the span of the initial ensemble.
We choose to use a Gaussian distribution due to its relationship with a squared-exponential kernel (Rahimi and Recht, 2007) and find that performance of the algorithm depends strongly on the choice of a covariance structure. Desiring a kernel that is flexible yet scales less quickly than the naive $\mathcal{O}((d+p)^{2})$ parameters of a Cholesky factored covariance, the low-rank perturbationstructure of Section 3.4 was chosen across all applications:
$\varphi(x;(\sigma,U,S))$ $=\sqrt{\sigma}\cos(\Xi x+B),\quad\text{where}\quad$
@@
To enforce positivity of $\sigma$ and the entries on the diagonal matrix $S_{ii}$, we assume independent Lognormal distributions with 99% of prior support covering the interval $(10^{-3},10^{3})$. For the entries $U_{ij}$, we took independent Gaussian prior distributions with 99% of the prior support covering $(-300,300)$. Over the different experiments, the dependence on these prior ranges was quite weak. They were observed however to influence speed of algorithm convergence when the prior span was overly concentrated or spread by several orders of magnitude.
@@ A.2 Common EKI Configuration
Besides the prior, we also detail the common EKI configuration parameters of the applications. In the experiments we chose to use the adaptive learning rate scheduler from (Iglesias and Yang, 2021). The adaptive value of $\Delta t_{n+1}$ provides a large step while bounding Jeffrey’s divergence between the distributions represented by the ensemble at $t_{n}$ to $t_{n+1}$. This prevents accumulation of errors due to ill-conditioned updates with large $\Delta t$. A small amount of additive white noise (inflation) was added to the particles at each iteration to prevent degeneracy in the ensemble exploration.
The input and output data was normalized before training. In particular, a whitening matrix was applied so that (i) the input samples have mean zero and unit variance. The output data was also whitened so that the variability $\Gamma$ had identity covariance over the output data. The $\Gamma(u)$ in Subsection 3.3, was always estimated at a fixed value of $u=u^{\dagger}$ taken as the mean of the prior, and an optimal linear shrinkage estimator was used to ensure a well-conditioned matrix (Ledoit and Wolf, 2004).
@@
To train the hyperparameters in the Ishigami experiment, the ensemble size was taken to be 30, for the 13 parameters #$\#\{\sigma,\{U_{ij}\}_{ij},\{D_{ii}\}_{i}\}=(1,9,3)$. The number of features used while optimizing the hyperparameters was taken to be 150. The convergence is plotted over 20 iterations, and under 20 restarts of the algorithm with different random samplings from the prior in Figure 7.
Figure 7: Convergence of the error $N_{\mathrm{ens}}^{-1}\sum_{n=1}^{N_{\mathrm{ens}}}\|\Gamma^{-\frac{1}{2}}(y_{n}-\mathcal{G}(u_{n}))\|^{2}$ for $20$ different training experiments with scalar-valued random features. The initial value is normalized to one. For this experiment, convergence stagnates at around $8$ ensemble iterations.
To train the hyperparameters in the G-function experiment of dimension $d$, the ensemble size was taken to be 200 across all dimensional experiments. For the parameters of the rank-$r$ kernel, #$\#\{\sigma,\{U_{ij}\}_{ij},\{D_{ii}\}_{i}\}=(1,20r,r)$, we took rank $r=\min(10,d)$. The number of features used while optimizing the hyperparameters was taken to be 300. The optimization was iterated over 20 iterations with a fixed ensemble size of 200 over all experiments. We do not plot the convergence for all experiments for brevity, but report that the normalized loss was reduced to around $10^{-2}$ over the iterations.

### v4

@@ 5 Conclusion
Ensemble Kalman inversion (EKI) and its variants have proven to be easily implementable and versatile for efficiently solving optimization problems with stochastic objective functions. Such problems arise when tuning hyperparameters in randomized algorithms, and specifically in the model selection problem found in randomized machine learning approaches. This paper explored a concrete implementation applied to random feature regression for function emulation. The main developments in this work were threefold. First, by viewing random feature emulators as Gaussian processes, the paper used empirical Bayesian arguments to construct an objective function for hyperparameter estimation. Second, this objective function was adapted to an inverse problem framework compatible with EKI. Third, the work provided structured choices of feature distributions which, when tuned with EKI, produced emulators that performed comparably to Gaussian processes and displayed robustness to overfitting and feature selection.
5.1 Discussion
@@
Random features can offer satisfactory approximation with fewer structural assumptions and computational requirements than Gaussian processes. As illustrated in Section 4.3, vector-valued random feature emulators were able to produce qualitatively similar posterior distributions without relying on diagonalization in the output space—though the framework can take advantage of diagonal structure when known. As with Gaussian processes, the choice of feature map and distribution is important. In this paper, the random Fourier features with feature distributions parametrized by low-rank perturbations (Subsection 3.4) were effective across all test cases for a suitably large rank. Further investigation of sensible feature distribution selection is left to future work.
Performance of random feature regression depends on balancing expressiveness with efficiency in several areas. One must select an appropriate feature distribution family that is both flexible and parameter-efficient for large input and output dimensions. One must propose an optimizer and objective function in order to tune the free parameters of this family efficiently and robustly. The proposed feature distribution families in Subsection 3.4 were shown to be sufficiently flexible for a wide range of applications, but required a rank that had to be selected offline. The objective function (45) contains a data misfit that scales with $Np$ and the noise factor $\Gamma$ with $(Np)^{2}$, where $N$ is the size of the training data set and $p$ the output dimension. The proposed adjustment in this paper is to use a cross-validation-inspired partitioning and an offline estimation of a constant $\Gamma$ to reduce these costs. Such changes afforded a feasible online cost of the objective function in our experiments, but we expect that a better balance of accuracy with efficiency can be gained through more principled approximation of $\Gamma$. The proposed EKI optimizer relies on selection of an ensemble size and a prior distribution. A success of EKI is that the ensemble size and parameter dimension are in fact only weakly coupled and newer variants use sampling error correction methods to decouple them further (Vishny et al., 2024). Such scalability properties are inherited from the ensemble Kalman filtering literature, where assimilators must routinely estimate states of size $\mathcal{O}(10^{9})$ with an ensemble size of $\mathcal{O}(10)$. Sensible choice of the prior distribution, here taken to be weakly informative independent distributions in each dimension, is still an open question. Since the proposed framework learns the hyperparameters of the feature distribution, the number of random feature samples $M$ is a free parameter. In practice, it is often advantageous to take $M$ smaller for hyperparameter learning (where accuracy is less important) and take $M$ larger for prediction tasks (where accuracy is the goal). To explore the practical significance of such benefits, comparison with other fixed-feature approaches is an interesting direction for further work, as is elucidating the relationship to large scale hyperparameter transfer methods (Yang and Hu, 2020; Yang et al., 2022)
@@
Along related lines, a fruitful area for future development is the design of alternative objective functions for hyperparameter tuning that further improve optimizer performance and accelerate convergence. Objective functions not motivated from Bayesian arguments exist for kernel methods and are based on cross-validation (Owhadi and Yoo, 2019; Chen et al., 2021; Naslidnyk et al., 2023). Such objective functions have proven successful in derivative-based settings for deterministic kernels and could also be recast for random objective functions such as those required in the present derivative-free framework. The development of theoretical guarantees, like those in Chen et al. (2021), is also of great importance. Furthermore, since EKI solves an inverse problem, misfits appearing in the objective function need not be based on directly observed input-output pairs. This opens the door to tuning hyperparameters from indirect data. Learning from indirect data is useful in many applications where one cannot directly observe the outputs of the model. For example, embedding a machine learning model into the right hand side of a dynamical system will often require online training for stability and performance, yet the input and output of the model is not observable (Pahlavan et al., 2024; Wu et al., 2024).
5.2 Outlook
@@
Broadly, the present paper intends to act as a case study within a more general framework for learning distributions that are sampled in randomized algorithms. Besides the random features studied here, other machine learning tools that follow a randomization approach (Scardapane and Wang, 2017) include Nyström methods (Williams and Seeger, 2000; Sun et al., 2015; Meanti et al., 2022) and reservoir computers (Gauthier, 2018; Gauthier et al., 2021; Griffith et al., 2019). Common patterns to tune the hyperparameters of such methods often follow the existing literature on random features: employing manual calibration, fitting to fixed realizations of random variables, and restricting to a narrow choice of algorithm- and problem-specific hyperparameters. The proposed framework is a principled and automated alternate strategy. Moreover, these fields have witnessed other Bayesian optimization methods learn low-dimensional (e.g., fewer than $10$) hyperparameters effectively (Griffith et al., 2019; Racca and Magri, 2021; Antonik et al., 2023). Since EKI is highly scalable when compared with such Gaussian process-based optimizers, it may enable the learning of much higher-dimensional hyperparameters. We remark that recent work also suggests Bayesian optimization methods may be scaled to high input and output spaces (Hvarfner et al., 2024; Xu et al., 2024; Maddox et al., 2021).
A parallel literature uses derivative-free evolutionary algorithms to search for deterministic network architectures in a process termed neural evolution and claims superiority over grid search and random search (Bergstra and Bengio, 2012) over discrete domains (Real et al., 2017; Stanley et al., 2019; Xiao et al., 2020). Such algorithms are also applied at times to continuum hyperparameter domains in low dimensions. All such methods are susceptible to the curse of dimensionality, however. Future work that systematically compares the effectiveness of EKI against Bayesian optimizers, evolutionary algorithms, and random search would be of great value to the field and more decisively determine whether EKI is better suited for higher-dimensional hyperparameters. Altogether, the methodological framework that the present paper develops lays the groundwork for novel automated approaches to hyperparameter calibration on different scales than the existing literature, while simultaneously providing robust and user-friendly software.
@@
The code and examples for this work are available in three open-source software packages written in the Julia programming language. RandomFeatures.jl v0.3.4 contains a range of different random feature families, EnsembleKalmanProcesses.jl v2.0.1 contains a range of EKI variants and utilities for creating and solving inverse problems, and CalibrateEmulateSample.jl v0.6 contains the code for the application problems presented in this paper.
@@
ORAD is supported by Schmidt Sciences, LLC, the National Science Foundation (NSF) under award number AGS-1835860, and the Office of Naval Research (ONR) under award number N00014-23-1-2654. NHN acknowledges support from NSF award number DMS-2402036, the NSF Graduate Research Fellowship Program under award number DGE-1745301, the Amazon/Caltech AI4Science Fellowship, the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation), and the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under ONR award number N00014-22-1-2790. The computations presented in this paper were partially conducted on the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology. The authors are grateful to Daniel J. Gauthier for helpful comments on a previous version of the paper and to the two anonymous referees for their valuable feedback.
@@ Appendix A Additional Numerical Experiment Details
In this appendix, we add some additional notes about the experimental configurations to facilitate reproducibility. Appendix A.1 discusses hyperparameter priors for the chosen feature distribution parametrizations, while the settings for the EKI algorithm are collected in Appendix A.2. The applications in Section 4 were chosen because they require satisfactory performance off of training data. In this appendix, we supplement these scientific applications with auxiliary investigations into the heuristics of running the proposed hyperparameter tuning algorithm. Summarizing the additional results found in Appendices A.3, A.4, and A.5, we find that few features—fewer than training data set size—are sufficient to obtain high-performing tuned hyperparameter, and the structure and rank of the feature distributions are important but can be selected with using a held-out validation data set.
Finally, we note that all numerical experiments are reproducible at
https://github.com/CliMA/CalibrateEmulateSample.jl .
Users may interact with this code base as a playground to understand our EKI-based algorithm and how its various setup parameters influence performance in the context of emulator training. Documentation can be found here:
https://clima.github.io/CalibrateEmulateSample.jl/dev/ .
Readers should navigate to Examples $\to$ Emulator testing.
@@ A.1 Hyperparameter Priors
To apply EKI, prior probability distributions must be placed on learnable parameters. From a practical perspective, priors serve as a method of introducing information, such as enforcing bounds and setting the initial ensemble span and correlation structure. In particular, for linear problems, the standard EKI update will be confined to the span of the initial ensemble. Thus, prior specification is a crucial component part of the framework.
We choose to use a Gaussian prior distribution due to its relationship with a squared-exponential kernel (Rahimi and Recht, 2007) and find that performance of the algorithm depends strongly on the choice of the prior’s covariance structure. Desiring a feature distribution that is flexible yet scales better than the naïve $\mathcal{O}((d+p)^{2})$ parameters of a Cholesky factored covariance, the low-rank perturbation structure from Section 3.4 was used in all applications:
$\varphi(x;(\varsigma,U,S))$ $=\sqrt{\varsigma}\cos(\Xi x+B),\quad\text{where}\quad$
@@
We also compare this class of nonseparable feature distributions with separable feature distributions in multi-output experiments. The class of separable distributions we chose factors both the input and output space, leading to
$\varphi(x;(\varsigma,V_{1},T_{1},V_{2},T_{2}))$ $=\sqrt{\varsigma}\cos(\Xi x+B),\quad\text{where}\quad$
$(\Xi,B)$ $\sim\mathrm{MatrixNormal}\bigl{(}0,C_{\mathrm{in}},C_{\mathrm{out}}\bigr{)}\otimes\mathsf{Unif}([0,2\pi]^{p})\,,$
$C_{\mathrm{in}}$ $\coloneqq(I+V_{1}T_{1}V_{1}^{\top})(I+V_{1}T_{1}V_{1}^{\top})^{\top}\,,\quad\text{and}\quad$
$C_{\mathrm{out}}$ $\coloneqq(I+V_{2}T_{2}V_{2}^{\top})(I+V_{2}T_{2}V_{2}^{\top})^{\top}.$
This separable kernel places a covariance structure on inputs (subscript $1$) and outputs (subscript $2$) independently and is more restrictive than the nonseparable distribution. However, the separable case has fewer hyperparameters. In the preceding displays, the matrices $S$, $T_{1}$, and $T_{2}$ are all diagonal.
To enforce positivity of $\varsigma$ and the entries of the diagonal matrices, we endow them with independent log-normal prior distributions with $99\%$ of prior support covering the interval $(10^{-3},10^{3})$. For the entries $U_{ij}$, we took independent Gaussian prior distributions with $99\%$ of the prior support covering $(-300,300)$. In the different experiments, the dependence on these prior ranges was quite weak. They were observed however to influence the speed of algorithm convergence when the prior span was overly concentrated or overly spread by several orders of magnitude.
@@ A.2 Common EKI Configuration
Besides the prior, we also detail the common EKI configuration parameters used across all three application experiments. In the experiments, we chose to use the adaptive learning rate scheduler from the work of Iglesias and Yang (2021). The adaptive value of step size $\Delta t_{n+1}$ provides a large step when bounding Jeffrey’s divergence between the distributions represented by the ensemble at times $t_{n}$ and $t_{n+1}$. This prevents accumulation of errors due to ill-conditioned updates with large $\Delta t$. A small amount of additive white noise (i.e., inflation) was added to the particles at each iteration to prevent degeneracy in the ensemble exploration. Additionally, the optimizer was more robust under multiple validation partitions $\Omega^{\textsf{c}}$ as described in Subsection 3.3. We chose to stack two partitions of the data during training and took $\mathsf{K}/N=0.2$ for all experiments except in Subsection 4.3, where we set $\mathsf{K}/N=0.1$. The input and output data was normalized before training. In particular, a whitening matrix was applied so that the input samples had empirical mean zero and unit empirical variance. The $\Gamma=\Gamma(u)$ in Subsection 3.3 was always estimated at a fixed value of $u=u^{\dagger}$ taken as the mean of the prior. The output data was also whitened so that the variability matrix $\Gamma$ had identity covariance over the output data. An optimal linear shrinkage estimator was used to ensure a well-conditioned matrix (Ledoit and Wolf, 2004).
@@
To tune the hyperparameters in the Ishigami experiment, the ensemble size was taken to be $J=30$ for the 13 parameters #$\#\{\varsigma,\{U_{ij}\}_{ij},\{D_{ii}\}_{i}\}=(1,9,3)$. The number of features used while optimizing the hyperparameters was taken to be 150. In Figure 7, the convergence is plotted over 20 iterations and under 20 restarts of the algorithm with different random samplings from the prior.
Figure 7: Convergence of the error $n_{\mathrm{iter}}^{-1}\sum_{n=1}^{n_{\mathrm{iter}}}\|\Gamma^{-1/2}(z_{n}-\mathcal{G}(u_{n}))\|^{2}$ for $20$ different tuning experiments with scalar-valued random features used to learn the Ishigami function for global sensitivity analysis. The initial objective value is normalized to one in the plot. For this experiment, convergence stagnates at around eight ensemble iterations.
To tune the hyperparameters in the G-function experiment of dimension $d$, the ensemble size was taken to be $200$ across all dimensional experiments. For the parameters of the rank-$r$ feature distribution, #$\#\{\varsigma,\{U_{ij}\}_{ij},\{S_{ii}\}_{i}\}=(1,20r,r)$, we took rank $r=\min(10,d)$. The number of features used while optimizing the hyperparameters was taken to be $M=300$. The optimization was run for 20 iterations with a fixed ensemble size of 200 over all experiments. We do not plot the convergence for all experiments for brevity, but report that the normalized objective value was reduced to under $10^{-1}$ over the iterations.
Table 3 displays timings of the tuning and fitting process for the Sobol G-Function against different input dimensions. Both GP and RF methods are compared. Since the algorithms are not reaching the same convergence criteria and performance may depend heavily on implementation details of numerical linear algebra subroutines under-the-hood, such timings are qualitative. However, these results provide evidence of the poor scaling of GP hyperparameter tuning even in moderate dimensions and even with a fixed number of input data points. On the other hand, EKI scales linearly and is also able to take advantage of multithreading—though some overheads are seen here.
Table 3: Mean time to train (tune and fit) the machine learning algorithms, averaged over $30$ repeated experiments for learning the Sobol G-function for global sensitivity analysis. Note this comparison runs the default convergence criteria for L-BFGS-B from SciKitLearn.jl and runs EKI for $20$ iterations using $200$ ensemble members. The final GP training was canceled after 10 hours. The memory for both tools in serial was 2GB, while EKI in parallel used approximately 6GB.
$d$ GP (L-BFGS-B) RF (EKI) – Serial RF (EKI) – $8$ Threads
3 16s 46s 12s
6 309s 65s 23s
10 2653s 134s 38s
20 $>$10h 270s 67s
A.3.1 Performance Dependence on Number of Features
A useful perk that the proposed hyperparameter tuning framework for RFs can utilize is a decoupling of the number of features $M$ required during tuning from the number of features $M^{\prime}\neq M$ required during prediction. The timings and performance of the algorithm for the six-dimensional Sobol G-function over varying numbers of features $M$ used in tuning is displayed in Table 4. The error here represents the $L^{2}$ difference between the predicted RF mean (with $M^{\prime}=1000$ features) and the noiseless data, averaged either over the fit inputs (training error) or the other sampling points (test error). In either case, it is observed that increasing the number of features leads to a scaling in time between linear and quadratic, and a reduction of both the training and test error. However, there is clear stagnation in performance with respect to both training and test error when the algorithm exceeds $M=100$ features. This is potentially due to sample error domination or optimization effects.
Table 4: The tuned performance for the six-dimensional Sobol G-function emulators as different numbers of features $M$ are taken during tuning. The training and test errors are the average $L^{2}$ difference between the tuned RF mean (created always with $M^{\prime}=1000$ features) and noiseless data at training and test points. The final row corresponds to the GP emulator. The time taken for the tuning process is also shown.
$M$ (When Tuning) Optimizer Time (Serial) Training Error Test Error
25 7s 0.014420 0.004687
50 13s 0.011280 0.004069
100 30s 0.007872 0.003643
200 72s 0.007720 0.003797
400 204s 0.006297 0.003535
800 700s 0.005756 0.003660
$\infty$ (GP Limit) 320s 0.000358 0.002827